OpenStudy (anonymous):

Find the direct variation equation which passes through (0, 0) and (4, 1)

ganeshie8 (ganeshie8):

direct variation : y = kx here, k is the slope.

ganeshie8 (ganeshie8):

you know the slope formula when two points are given ?

OpenStudy (anonymous):

not really

ganeshie8 (ganeshie8):

okay, slope between (x1, y1) and (x2, y2) = \(\huge \frac{y2-y1}{x2-x1}\) k = slope between ( 0, 0) and ( 4, 1) = \(\huge \frac{1-0}{4-0} = \frac{1}{4}\) just plugin "k" in : y = kx the direct variation equation becomes, \(\huge y = \frac{1}{4}x\)

ganeshie8 (ganeshie8):

@DNADRE does it help ?

OpenStudy (anonymous):

yea

ganeshie8 (ganeshie8):

glad to hear :)